# Parafermions in Moiré Minibands

### v1

@@ Abstract
Moiré materials provide a remarkably tunable platform for topological and strongly correlated quantum phases of matter. Very recently, the first zero field Abelian fractional Chern insulators (FCIs) have been experimentally demonstrated and it has been theoretically predicted that non-Abelian states with Majorana fermion excitations may be realized in the nearly dispersionless minibands of these systems. Here we provide telltale evidence in terms of low-energy quantum numbers, spectral flow and entanglement spectra for the even more exotic possibility of moiré-based non-Abelian FCIs exhibiting Fibonacci parafermion excitations.
@@
Introduction.—In the last years, moiré materials have been established as an accessible and highly tunable platform for exploring strongly correlated topological phases of matter [1, 2, 3, 4, 5, 6, 7]. Most notably, specific van der Waals heterostructures consisting of twisted layers of graphene or transition metal dichalcogenides have been theoretically predicted [8, 9, 10, 11, 12, 13, 14, 15] and experimentally observed [16, 17, 18, 19, 20, 21, 22] to host fractional Chern insulators (FCIs)—lattice analogs of fractional quantum Hall (FQH) states that can exist in the absence of a magnetic field [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. So far, the FCI states that have been unambiguously identified in experiments are topologically equivalent to Abelian hierachical FQH states [34, 35, 36, 37] hosting Abelian anyon excitations [38, 39]. While the possibility of exploring anyon physics at zero magnetic field and moderate temperatures is already remarkable, the pursuit of non-Abelian quasiparticles—which should appear as elementary excitations in more exotic FCI states—constitutes a more ambitious and potentially rewarding venture with prospects for fault-tolerant topological quantum computation [40].
@@
Recent experimental signatures [41], suggested to arise from non-Abelian topological order at moiré filling fractions with even denominator, have sparked a series of theoretical works addressing the stability of Moore-Read (MR) states in moiré systems [42, 43, 44, 45, 46, 47, 48, 49]. Strengthening the theoretical evidence based on the ground state degeneracy, which can correspond to either the MR phase or charge density waves (CDWs), further analysis counting quasi-hole excitations in entanglement spectra has shown that MR states could indeed be stable in moiré materials [47, 42, 48]. Unfortunately, the braiding operations allowed by the emergent Majorana excitations in MR states cannot generate any arbitrary unitary transformation and would thus be insufficient for universal quantum computation [40]. This issue can, however, be overcome by exploiting parafermions with richer braiding statistics, which have been predicted to appear in FQH systems [50], FQH–superconductor heterostructures [51], and certain non-Hermitian systems [52], although their experimental realization remains elusive. A natural step forward in the context of moiré FCIs is thus to search for Fibonacci parafermions appearing in the $\mathbb{Z}_{3}$ Read-Rezayi (RR) state [50]—which is characterized by the clustering of composite fermions into triplets, as opposed to the pairing in the MR phase.
@@
In this Letter, we provide conclusive evidence showing that the non-Abelian RR FCI at 3/5 filling can be realized in a moiré system. In particular, we perform many-body exact diagonalization on double twisted bilayer graphene (dTBG) and obtain the expected number of degenerate ground states appearing at momenta fulfilling the generalized Haldane statistics. Furthermore, we verify the persistence of the many-body energy gap upon insertion of an external magnetic flux, and, more strikingly, we show that the state counting in the low-energy sector of the particle-cut entanglement spectrum (PES) exactly matches the number of allowed quasi-hole excitations. In addition, we find that the RR phase is destroyed when the average quantum metric of the considered flat band approaches asymptotically the ideal value for the second Landau level (LL), highlighting both the usefulness and the limitations of this quantity as a heuristic indicator for the emergence of non-Abelian phases. Finally, we show that the RR phase is less stable at $2/5$ filling but still present—as reflected in the quasi-particle excitations accessed through the hole entanglement spectrum—, signalling the robustness of moiré-based parafermion FCIs.
@@
Figure 1: Evidence for moiré parafermions at $\nu=3/5$. (a) Low-lying many-body energy spectrum, (b) spectral flow, and (c) $N_{A}=5$ particle-cut entanglement spectrum for a $N_{\text{s}}=20$-site system at band filling $\nu=3/5$ and $\gamma=3.75$. The corresponding results for $N_{\text{s}}=25$ sites are shown in (d)-(f). The 10-fold nearly-degenerate ground states are marked by red dots. In the PES, the number of states below the red solid line is $14404$ for $N_{\text{s}}=20$ and $51255$ for $N_{\text{s}}=25$, both matching the quasi-hole excitation counting for Read-Rezayi states detailed in the SM. (g) Structure factor and (h) pair-correlation function of the Read-Rezayi ground states for $N_{\text{s}}=25$. The black lines in (h) denote the unit cell of the system. The spanning vectors for the considered systems are $\mathbf{T}_{1}=(3,4)$ and $\mathbf{T}_{2}=(-2,4)$ for $N_{\text{s}}=20$, and $\mathbf{T}_{1}=(5,0)$ and $\mathbf{T}_{2}=(0,5)$ for $N_{\text{s}}=25$.
@@
Setup.—We consider dTBG as an exemplary moiré system with the valley- and spin-polarized Hamiltonian [53]
$H_{0}(\bf r)=\begin{pmatrix}u_{1}I&D_{1}^{\dagger}(\bf r)&0&0\\ D_{1}(\bf r)&u_{2}I&\gamma I&0\\ 0&\gamma I&u_{3}I&D_{2}^{\dagger}(\bf r)\\ 0&0&D_{2}(\bf r)&u_{4}I~{}\end{pmatrix},$ (1)
[…]
In order to tackle the many-body problem of interacting electrons, we diagonalize $H_{0}(\bf r)$ in the basis of Bloch states and project the electron-electron interactions onto the considered nearly-flat band, obtaining $H_{\text{int}}=\frac{1}{2}\sum_{\bf k_{1}k_{2}k_{3}k_{4}}V_{\bf k_{1}k_{2}k_{3}k_{4}}c_{\bf k_{1}}^{\dagger}c_{\bf k_{2}}^{\dagger}c_{\bf k_{3}}c_{\bf k_{4}}.$ Here, the matrix element $V_{\bf k_{1}k_{2}k_{3}k_{4}}$ ensures momentum conservation and contains single-particle form factors [60, 8] as well as the 2D bare Coulomb potential $V({\bf q})=~{}\frac{1}{2A\epsilon|{\bf q}|}$ with the average dielectric constant of the system $\epsilon=10$ [61]. Next, we obtain the low-lying energy spectrum of many-body states by exact diagonalization of the spin- and valley-polarized interaction Hamiltonian $H_{\text{int}}$. In particular, we consider a periodic system of $N_{\text{e}}$ electrons in $N_{\text{s}}$ moiré unit cells with the filling $\nu=N_{\text{e}}/N_{\text{s}}=3/5,2/5$.
@@
Read-Rezayi parafermion states.—We start by considering a fractionally filled dTBG with $\nu=3/5$ and a relatively large coupling $\gamma=3.75$ between the two TBG sheets. Performing exact diagonalization on systems with size $N_{\text{s}}=20$ and $N_{\text{s}}=25$, we obtain 10 nearly-degenerate ground states separated from the excited states by an energy gap, cf. red dots in Fig. 1(a),(d). The splitting of these states is generically expected on the relatively small finite size systems accessible to exact diagnalization, but should go away in the large system limit. Moreover, we find that the ground states do not flow into excited states upon insertion of a magnetic flux, confirming the presence of a many-body gap, cf. Fig. 1(b),(e). The 10-fold ground state degeneracy is in agreement with that expected from RR states in the thin-torus limit, where the ground states at filling $\nu=k/(kM+2)$ are formed by $k$ particles occupying $kM+2$ consecutive sites and every two particles being separated by at least $M$ sites [62, 50, 63]. For fermionic Read-Rezayi states with $k=3$ and $M=1$, i.e. $\nu=3/5$, two distinct configurations are then expected, namely $1110011100\cdots$ and $1101011010\cdots$, which together with their translation-invariant partners result in a 10-fold degenerate ground state. Furthermore, the calculated ground states in Fig. 1(a),(d) appear at center-of-mass momenta matching those expected from the aforementioned thin-torus configurations.
@@
While the approximate ground state degeneracy and the spectral flow is consistent with the RR phase, it could also correspond to weakly entangled states such as CDWs—as has been shown for candidate MR states [47]. Thus, more information is needed to conclusively determine whether the ground states in Fig. 1(a),(d) correspond to RR states. To this end, we compute the PES of the ground states [64, 23]. The PES is obtained by dividing the many-body system into $A$ and $B$ subsystems consisting of $N_{A}$ and $N_{B}=N_{\text{e}}-N_{A}$ particles, and then calculating the eigenvalues of $-\mathrm{log}\rho_{A}$, where $\rho_{A}=\mathrm{tr}[~{}\frac{1}{N_{\text{d}}}~{}\sum_{i=1}^{N_{\text{d}}}~{}\ket{\Psi_{i}}\bra{\Psi_{i}}]$ is the reduced density matrix of $A$. Here, $\rho_{A}$ carries crucial information about the quasi-hole excitations in the $N_{\text{d}}$-fold degenerate ground states |$\ket{\Psi_{i}}$, which are characteristically distinct for Abelian FCIs, non-Abelian FCIs, and CDWs. In particular, a gap in the PES is expected, below which the number of states exactly matches the amount of quasi-hole excitations allowed by the specific quantum phase of the system.
@@
In Fig. 1(c),(f) we show the PES with $N_{A}=5$ for the two considered system sizes. In addition, we mark with a red line the entanglement energy below which the number of states matches exactly the amount of allowed quasi-hole excitations (the analytical counting is detailed in the SM [55]). For the smaller system ($N_{\text{s}}=20$), the dense low-energy sector is followed by a region of sparsely distributed states resembling a gap (around the red line in Fig. 1(c)), below which the state counting is similar to that expected for the RR phase. Remarkably, the PES for the larger system ($N_{\text{s}}=25$) shows a clear entanglement gap with the low-energy state counting being exactly equal to the number of quasi-hole excitations in the RR phase. This observation serves as smoking-gun evidence that the considered system is in the RR phase.
@@
To further clarify the nature of the phase, we calculate the structure factor and pair-correlation function averaged over the tenfold quasi-degenerate ground states, which provide insights on the crystalline or liquid character of the system. Here, the structure factor helps measure if the ground states exhibit a preferred charge order and is defined as $S(\mathbf{q})=\frac{1}{N_{\text{e}}}(\langle\rho(\mathbf{q})\rho(-\mathbf{q})\rangle-\delta_{\mathbf{q},0}N_{\text{e}}\langle\rho(\mathbf{q})\rangle),$ with $\rho(\mathbf{q})$ being the density operator projected onto the nearly flat band. On the other hand, the real space pair-correlation function reads $g(\mathbf{r})=~{}\frac{1}{N^{2}_{\text{e}}}(\braket{n(\mathbf{r})n(\mathbf{0})}-\delta_{\mathbf{r},0}\braket{n^{2}(\mathbf{r})})$, where $n(\mathbf{r})$ is the real-space density operator, and can be obtained via the Fourier transform of $S(\mathbf{q})$. For liquid phases, $g(\mathbf{r})$ approaches a constant for large $\mathbf{r}$, while it shows a periodic arrangement of charges for CDWs. As shown in Fig. 1(g)-(h), both the structure factor and the pair-correlation function indicate a liquid behavior—concretely, no prominent peaks in $S(\mathbf{q})$ and no CDW unit cells appear in $g(\mathbf{r})$.
@@
Figure 2: Stability of the parafermion phase and quantum geometry. The many-body energy spectrum gap for non-Abelian $\mathbb{Z}_{3}$ RR states at $\nu=3/5$ (red dotted lines, right axis) as a function of the coupling strength $\gamma$ between twisted layers is shown for system sizes $N_{\text{s}}=20$ (squares) and $N_{\text{s}}=25$ (circles) together with the averaged quantum geometry (blue line, left axis).
@@
Stability of parafermion FCIs and quantum geometry.—Having demonstrated the existence of non-Abelian RR states, we now explore the stability of such topological phases in the parameter space associated with the quantum geometry, which serves as a heuristic indicator of the analogy between the partially filled moiré band and the LLs in two-dimensional electron gases [49]. For the $n$-th LL, the quantum geometry and the Berry curvature satisfy [57] $\text{tr}[g(\mathbf{k})]=(2n+1)|\Omega(\mathbf{k})|$, with $n=0,1,2...$. It is then generally believed that a moiré band is expected to mimic the respective LL physics if this relation is approximately satisfied. Following this principle, it has been recently shown that both the Laughlin-like zero-energy modes [10] and non-Abelian MR state can be realized in a moiré flat band with a (nearly) ideal quantum geometry [42, 43, 44, 45, 46, 48]. As mentioned before, the average quantum geometry $\chi$ of the considered flat band in dTBG increases gradually from $1$ to $3$ as $\gamma$ is increased from 00 to infinity (see blue line in Fig. 2). The parameter $\gamma$ then provides an ideal tuning knob to study the transition between the physics of the lowest LL and the second LL.
@@
In the strong coupling limit, $\gamma\rightarrow\infty$, the targeted band should in principle emulate second LL physics, as it acquires an ideal quantum geometry $\chi=3$ and the band dispersion becomes exactly flat. However, while the RR states are energetically competitive in the first excited Landau level [58, 59], the many-body low-lying energy spectrum gap shown in Fig. 2 indicates their absence as $\gamma\rightarrow\infty$. Instead, the $\mathbb{Z}_{3}$ RR appears at an intermediate coupling strength, ranging from $\gamma\approx 3$ to $4.5$. Here, the energy gap is defined as $\Delta_{E}=E_{\text{e}}-E_{\text{g}}$ with $E_{\text{e}}$ being the $11$th lowest energy and $E_{\text{g}}$ being the $10$th lowest energy. These findings strongly suggest that the quantum geometry criterion is not sufficient to diagnose the non-Abelian topological phases. It nevertheless gives some intuition of where exotic states may occur—the RR states do in fact form in a region where the quantum geometry resembles that of the first excited rather than that of the lowest LL. We note that there are alternative ways of making moiré minibands that are close but not identical to the first excited LL such as considering other magic angles [65], adding a “negative” magnetic field [66] or periodic strain [67], and considering bands further away from charge neutrality [42, 46, 48, 43].
@@
In the weak coupling limit $\gamma\rightarrow 0$, where the system approaches to two decoupled chiral TBG sheets, the average quantum geometry $\chi\rightarrow 1$ shown in Fig. 2 suggests that the targeted flat band resembles the lowest LL, where Abelian hierarchy states at band filling $\nu=3/5$ and its particle-hole partner $\nu=2/5$ [35, 36, 37] are indeed present (see the SM [55] for more details).
@@
Particle-hole asymmetry and hole entanglement spectrum.—We now search for the possibility of RR states at electronic band filling $\nu=2/5$, which in ideal LLs correspond to the particle-hole conjugate of the $\nu=3/5$ RR phase. It has been shown that, in general, particle-hole symmetry is broken in FCIs and moiré systems [68, 8], a fact that can bee seen as a consequence of a fluctuating quantum geometry  [69, 70]. In line with this asymmetry, we do not find a clear many-body energy gap of the RR states for filling $\nu=2/5$ at the system sizes considered, suggesting that the RR phase is less stable for $\nu=2/5$ than for $\nu=3/5$ (see Fig. 3(a)). In addition, although the tenfold quasi-degenerate states are still present at the correct momentum sector, their corresponding PES does not show a gap with the correct quasi-hole counting (see the SM [55]). This, however, should not be surprising: since the $\nu=2/5$ RR phase can be understood as the RR phase for holes at 3/5 hole-filling, adding holes to the system directly increases the ground-state energy as the 4-body interaction can no longer be minimized.
@@
Instead, the $\nu=2/5$ RR phase allows a certain number of quasi-particle excitations without increasing the ground-state energy. This is reflected in the hole entanglement spectrum (HES), where the many-body system is divided into $A$ and $B$ subsystems with $N_{A}$ and $N_{B}=N_{\text{s}}-N_{\text{e}}-N_{A}$ now denoting the number of holes. Importantly, the HES (Fig. 3(b)) displays an entanglement gap below which the number of states matches the quasi-particle counting of the 3/5 hole-filling (i.e. $\nu=2/5$) RR phase, indicating that the calculated states indeed correspond to the non-Abelian RR phase. From a topological perspective, this then implies that the nature of the RR states remains largely unaffected by the renormalization of the hole dispersion that is expected to arise from fluctuations in the quantum geometry [68], even though these fluctuations may alter the ground state energies as suggested by the absence of a gap in Fig. 3(a).

### v2

@@ Abstract
Moiré materials provide a remarkably tunable platform for topological and strongly correlated quantum phases of matter. Very recently, the first Abelian fractional Chern insulators (FCIs) at zero magnetic field have been experimentally demonstrated, and it has been theoretically predicted that non-Abelian states with Majorana fermion excitations may be realized in the nearly dispersionless minibands of these systems. Here, we provide telltale evidence based on many-body exact diagonalization for the even more exotic possibility of moiré-based non-Abelian FCIs exhibiting Fibonacci parafermion excitations. In particular, we obtain low-energy quantum numbers, spectral flow, many-body Chern numbers, and entanglement spectra consistent with the $\mathbb{Z}_{3}$ Read–Rezayi parafermion phase in an exemplary moiré system with tunable quantum geometry. Our results hint towards the robustness of moiré-based parafermions and encourage the pursuit in moiré systems of these non-Abelian quasiparticles that are superior candidates for topological quantum computing.
Introduction
@@
In the last years, moiré materials have been established as an accessible and highly tunable platform for exploring strongly correlated topological phases of matter [1, 2, 3, 4, 5, 6, 7]. Most notably, specific van der Waals heterostructures consisting of twisted layers of graphene or transition metal dichalcogenides have been theoretically predicted [8, 9, 10, 11, 12, 13, 14, 15] and experimentally observed [16, 17, 18, 19, 20, 21, 22] to host fractional Chern insulators (FCIs)—lattice analogs of fractional quantum Hall (FQH) states that can exist in the absence of a magnetic field [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. So far, the FCI states that have been unambiguously identified in experiments are topologically equivalent to Abelian hierachical FQH states [34, 35, 36, 37] hosting Abelian anyon excitations [38, 39]. While the possibility of exploring anyon physics at zero magnetic field and moderate temperatures is already remarkable, the pursuit of non-Abelian quasiparticles—which should appear as elementary excitations in more exotic FCI states—constitutes a more ambitious and potentially rewarding venture with prospects for fault-tolerant topological quantum computation [40].
@@
Recent experimental signatures [41], suggested to arise from non-Abelian topological order at moiré filling fractions with even denominator, have sparked a series of theoretical works addressing the stability of Moore–Read (MR) states in moiré systems [42, 43, 44, 45, 46, 47, 48, 49]. Strengthening the theoretical evidence based on the ground state degeneracy, which can correspond to either the MR phase or charge density waves (CDWs), further analysis counting quasi-hole excitations in entanglement spectra has shown that MR states could indeed be stable in moiré materials [47, 42, 48]. Unfortunately, the braiding operations allowed by the emergent Majorana excitations in MR states cannot generate any arbitrary unitary transformation and would thus be insufficient for universal quantum computation [40]. This issue can, however, be overcome by exploiting parafermions with richer braiding statistics, which have been predicted to appear in FQH systems [50], FQH–superconductor heterostructures [51, 52], and certain non-Hermitian systems [53], although their experimental realization remains elusive. A natural step forward in the context of moiré FCIs is thus to search for Fibonacci parafermions appearing in the $\mathbb{Z}_{3}$ Read–Rezayi (RR) state [50]—which is characterized by the clustering of composite fermions into triplets, as opposed to the pairing in the MR phase. Realizing parafermion excitations in the widely accessible and tunable moiré materials would open an exceptionally promising avenue in the development of fault-tolerant topological quantum computers.
@@
In this work, we provide numerical evidence showing that the non-Abelian RR FCI at 3/5 filling can be realized in a moiré system. In particular, we perform many-body exact diagonalization on an exemplary moiré flat band based on a double twisted bilayer graphene (dTBG) model with tunable quantum geometry. First, we obtain the expected number of degenerate ground states appearing at momenta fulfilling the generalized Haldane statistics. After that, we verify the persistence of the many-body energy gap upon insertion of magnetic flux (i.e. twisted boundary conditions) and find that, on average, each ground state has a many-body Chern number of $3/5$. More strikingly, we show that the state counting in the low-energy sector of the particle-cut entanglement spectrum (PES) exactly matches the number of allowed quasi-hole excitations. In addition, we find that the RR phase is destroyed when the average quantum metric of the considered flat band approaches asymptotically the ideal value for the second Landau level (LL), highlighting both the usefulness and the limitations of this quantity as a heuristic indicator for the emergence of non-Abelian phases. Finally, we show that the RR phase is less stable at $2/5$ filling but still present—as reflected in the quasi-particle excitations accessed through the hole entanglement spectrum—, signaling the robustness of moiré-based parafermion FCIs.
@@
Figure 1: Evidence for moiré parafermions at $\nu=3/5$. (a) Low-lying many-body energy spectrum, (b) spectral flow, and (c) $N_{A}=5$ particle-cut entanglement spectrum for a $N_{\text{s}}=20$-site system with $\gamma=3.75$. The corresponding results for $N_{\text{s}}=25$ sites are shown in (d)-(f). The 10-fold nearly-degenerate ground states are marked by red dots. In the PES, the number of states below the red solid line is $14404$ for $N_{\text{s}}=20$ and $51255$ for $N_{\text{s}}=25$, both matching the quasi-hole excitation counting for Read–Rezayi states detailed in Supplementary Note 2. (g) Structure factor in the mini Brillouin zone and (h) pair-correlation function in the real-space unit cell for the Read–Rezayi ground states with $N_{\text{s}}=25$. The spanning vectors for the considered systems are $\mathbf{T}_{1}=(3,4)$ and $\mathbf{T}_{2}=(-2,4)$ for $N_{\text{s}}=20$, and $\mathbf{T}_{1}=(5,0)$ and $\mathbf{T}_{2}=(0,5)$ for $N_{\text{s}}=25$.
Results
@@
Setup. We consider an exemplary model based on dTBG describing moiré minibands with tunable quantum geometry. Concretely, the single-particle valley- and spin-polarized Hamiltonian reads [54]
$H_{0}(\bf r)=\begin{pmatrix}u_{1}I&D^{\dagger}(\bf r)&0&0\\ D(\bf r)&u_{2}I&\gamma I&0\\ 0&\gamma I&u_{3}I&D^{\dagger}(\bf r)\\ 0&0&D(\bf r)&u_{4}I~{}\end{pmatrix},$ (1)
where the 2x2 matrix $D_{11}({\bf r})=D_{22}({\bf r})=-2ik_{\theta}^{-1}\bar{\partial}$, $D_{12}({\bf r})=D_{21}(-{\bf r})=\alpha U({\bf r})$ describes TBG in the chiral limit [55], $\gamma$ is the coupling between the two TBG sheets, and $u_{i}$ is a layer/sublattice potential. Here, $k_{\theta}=4\pi/3a_{m}$ ($a_{m}$ is the moiré lattice constant), ¯$\bar{\partial}=\frac{1}{2}(\partial_{x}+i\partial_{y})$, and $U(\mathbf{r})=e^{-i\mathbf{q}_{1}\cdot\bf r}+e^{i\phi}e^{-i\mathbf{q}_{2}\cdot\bf r}+e^{-i\phi}e^{-i\mathbf{q}_{3}\cdot\bf r}$ with $\phi=2\pi/3$ and $\mathbf{q}_{n}=C_{3}^{n}\mathbf{k}_{\theta}$. Throughout this work, we consider the first magic angle [55], i.e. $\alpha\approx 0.586$, and set $u_{1}=0.1,u_{2}=u_{3}=u_{4}=0$ in order to obtain an isolated flat band.
In this setting, the system is characterized by two consecutive non-degenerate and nearly flat bands exhibiting equal Chern number $\mathcal{C}=1$ and becoming degenerate in the limit $\gamma\to\infty$ (see Supplementary Note 1 and Supplementary Figure 1). The upper band maintains an ideal quantum geometry [56] $\text{tr}[g(\bf k)]=|\Omega(\bf k)|$ as $\gamma$ varies, where $g(\bf k)$ and $\Omega(\bf k)$ are the Fubini–Study (FS) metric and the Berry curvature, respectively. A definition of the FS metric is provided in the Methods section. We focus on the lower band, which has a tunable FS metric with the average $\chi=\frac{1}{2\pi}\int_{\text{BZ}}d^{2}{\bf k}\,~{}\text{tr}[g(\bf k)]$ ranging from $\chi=1$ as $\gamma~{}\to 0$ to $\chi=3$ as $\gamma~{}\to~{}\infty$ corresponding to the average quantum metric of the lowest and the first excited (second) LLs, respectively [57]. We note, however, that this correspondence generally demands additional conditions related to vortexability [54]. Motivated by the fact that the RR states are energetically more competitive in the second LL of FQH systems [50, 58, 59], we employ $\gamma$ as a tuning knob to explore the emergence of this non-Abelian phase in our model.
@@
Read–Rezayi parafermion states. We start by considering a fractionally filled dTBG with $\nu=3/5$ and a relatively large coupling $\gamma=3.75$ between the two TBG sheets. Performing exact diagonalization on systems with $N_{\text{s}}=20$ and $N_{\text{s}}=25$ moiré sites (see Methods), we obtain 10 nearly-degenerate ground states separated from the excited states by an energy gap, cf. red dots in Fig. 1(a),(d). The splitting of these states is generically expected on the relatively small finite size systems accessible to exact diagonalization, but should go away in the large system limit. Moreover, we find that the ground states do not flow into excited states upon flux insertion (i.e. twisted boundary conditions), confirming the presence of a many-body gap, cf. Fig. 1(b),(e). The 10-fold ground state degeneracy is in agreement with that expected in the thin-torus limit, where a quantum Hall system is adiabatically mapped into a 1D lattice that can be exactly solved [60]. In particular, the RR ground states at filling $\nu=k/(kM+2)$ are formed by $k$ particles occupying $kM+2$ consecutive sites and every two particles being separated by at least $M$ sites [61, 50, 62]. Note that the 1D lattice sites in the thin torus are analogous to momentum points [23]. For fermionic Read–Rezayi states with $k=3$ and $M=1$, i.e. $\nu=3/5$, two distinct configurations are then expected, namely $1110011100\cdots$ and $1101011010\cdots$ (expressed in the occupation number representation), which together with their translation-invariant partners result in a 10-fold degenerate ground state. Furthermore, the calculated ground states in Fig. 1(a),(d) appear at center-of-mass momenta matching those expected from the aforementioned thin-torus configurations.
@@
While the approximate ground state degeneracy and the spectral flow are consistent with the $\mathbb{Z}_{3}$ RR phase, they could also correspond to weakly entangled states such as CDWs—as has been shown for candidate MR states [47]. Thus, more information is needed to conclusively determine whether the ground states in Fig. 1(a),(d) correspond to RR states. An additional aspect pointing towards the RR phase is the many-body Chern number, which we calculate to be $3/5$ for each ground state on average. We take one step further and obtain unambiguous evidence by computing the PES of the ground states [63, 23]. The PES, which is distinct than the entanglement entropy and provides much richer information, is obtained by dividing the many-body system into $A$ and $B$ subsystems consisting of $N_{A}$ and $N_{B}=N_{\text{e}}-N_{A}$ particles, and then calculating the eigenvalues of $-\mathrm{log}\rho_{A}$, where $\rho_{A}=\mathrm{tr}_{B}[~{}\frac{1}{N_{\text{d}}}~{}\sum_{i=1}^{N_{\text{d}}}~{}\ket{\Psi_{i}}\bra{\Psi_{i}}]$ is the reduced density matrix of $A$. Here, $\rho_{A}$ carries crucial information about the quasi-hole excitations in the $N_{\text{d}}$-fold degenerate ground states |$\ket{\Psi_{i}}$, which are characteristically distinct for Abelian FCIs, non-Abelian FCIs, and CDWs. In particular, a gap in the PES is expected, below which the number of states exactly matches the amount of quasi-hole excitations allowed by the specific quantum phase of the system.
@@
In Fig. 1(c),(f) we show the PES with $N_{A}=5$ for the two considered system sizes. In addition, we mark with a red line the entanglement energy below which the number of states matches exactly the amount of allowed quasi-hole excitations (the analytical counting is detailed in Supplementary Note 2). For the smaller system ($N_{\text{s}}=20$), the dense low-energy sector is followed by a region of sparsely distributed states resembling a gap (around the red line in Fig. 1(c)), below which the state counting is similar to that expected for the RR phase. Remarkably, the PES for the larger system ($N_{\text{s}}=25$) shows a clear entanglement gap with the low-energy state counting being exactly equal to the number of quasi-hole excitations in the RR phase. This observation serves as smoking-gun evidence that the considered system is in the $\mathbb{Z}_{3}$ RR phase.
@@
To further clarify the nature of the phase, we calculate the structure factor $S(\mathbf{q})$ and pair-correlation function $C(\mathbf{r})$ averaged over the tenfold quasi-degenerate ground states, which provide insights on the crystalline or liquid character of the system (see Methods). For liquid phases, $C(\mathbf{r})$ approaches a constant for large $\mathbf{r}$, while it shows a periodic arrangement of charges for CDWs. As shown in Fig. 1(g)-(h), both the structure factor and the pair-correlation function indicate a liquid behavior—concretely, no prominent symmetry-breaking peaks appear in $S(\mathbf{q})$ and $C(\mathbf{r})$ is constant.
@@
Figure 2: Stability of the parafermion phase and quantum geometry. (a) The many-body energy spectrum gap for non-Abelian $\mathbb{Z}_{3}$ RR states at $\nu=3/5$ (red dotted lines, right axis) as a function of the coupling strength $\gamma$ between twisted layers is shown for system sizes $N_{\text{s}}=20$ (squares) and $N_{\text{s}}=25$ (circles) together with the averaged quantum metric $\chi$ (blue line, left axis). The energy gap is defined as the difference between the energies of the lowest non-RR state and the highest RR state (typically the $11$th and the $10$th lowest energies), and it is set to zero if the difference is negative. The shaded area denotes the regime where parafermions are stable. (b) Quantum metric $g(\mathbf{k})A_{\text{BZ}}/2\pi$ of the single-particle band for $\gamma=0.20$ and (c) $\gamma=3.75$. $A_{\text{BZ}}$ is the Brillouin zone area.
@@
Stability of parafermion FCIs and quantum geometry. Having demonstrated the existence of non-Abelian RR states, we now explore the stability of such topological phases in the parameter space associated with the quantum geometry, which serves as a heuristic indicator of the analogy between the partially filled moiré band and the LLs in two-dimensional electron gases [49, 64, 65, 66, 67, 68], although we note again that an accurate correspondence generally demands additional conditions [54]. For the $n$-th LL, the quantum metric and the Berry curvature satisfy [57] $\text{tr}[g(\mathbf{k})]=(2n+1)|\Omega(\mathbf{k})|$, with $n=0,1,2...$. Motivated by this property of LLs, it has been recently shown that both the Laughlin-like zero-energy modes [10] and non-Abelian MR state can be realized in a moiré flat band with a (nearly) ideal quantum geometry [42, 43, 44, 45, 46, 48]. As mentioned before, the average quantum metric $\chi$ of the considered flat band in dTBG increases gradually from $1$ to $3$ as $\gamma$ is increased from 00 to infinity (see blue line in Fig. 2(a)). The parameter $\gamma$ then provides an ideal tuning knob to study the transition between the physics of bands with quantum geometric properties similar to the lowest LL and the second LL.
@@
In the strong coupling limit, $\gamma\rightarrow\infty$, the targeted band acquires an ideal average quantum metric $\chi=3$ and the band dispersion becomes exactly flat as in the second LL. However, while the RR states are energetically competitive in the first excited Landau level [58, 59], the many-body low-lying energy spectrum gap shown in Fig. 2(a) indicates their absence as $\gamma\rightarrow\infty$. Instead, the $\mathbb{Z}_{3}$ RR phase appears at an intermediate coupling strength, ranging from $\gamma\approx 3$ to $4.5$. We note that the fluctuation of the quantum metric is minimized around this range of $\gamma$ (cf. Fig. 2(b)-(c)) and increases slightly for larger $\gamma$ values (see Supplementary Figure 2). These findings strongly suggest that the quantum geometry criterion is not sufficient to diagnose the non-Abelian topological phases. It nevertheless gives some intuition of where exotic states may occur—the RR states do in fact form in a region where the quantum geometry resembles that of the first excited rather than that of the lowest LL. We note that there are alternative ways of making moiré minibands that are close but not identical to the first excited LL such as considering other magic angles [69], adding a “negative” magnetic field [70] or periodic strain [71], and considering bands further away from charge neutrality [42, 46, 48, 43].
@@
In the weak coupling limit $\gamma\rightarrow 0$, where the system approaches to two decoupled chiral TBG sheets, the average quantum metric $\chi\rightarrow 1$ shown in Fig. 2(a) and the quantum geometry distribution shown in Fig. 2(b) and Supplementary Figure 2 suggest that the targeted flat band resembles the lowest LL, where Abelian hierarchy states at band filling $\nu=3/5$ and its particle-hole partner $\nu=2/5$ [35, 36, 37] are indeed present (see Supplementary Note 3).
@@
Particle-hole asymmetry and hole entanglement spectrum. We now search for the possibility of RR states at electronic band filling $\nu=2/5$, which in ideal LLs correspond to the particle-hole conjugate of the $\nu=3/5$ RR phase. It has been shown that, in general, particle-hole symmetry is broken in FCIs and moiré systems [72, 8]—a fact that can be seen as a consequence of a fluctuating quantum metric [73, 74]. In line with this asymmetry, we do not find a clear many-body energy gap of the RR states for filling $\nu=2/5$ at the system sizes considered, suggesting that the RR phase is less stable for $\nu=2/5$ than for $\nu=3/5$ (see Fig. 3(a)). In addition, although the tenfold quasi-degenerate states are still present at the correct momentum sector, their corresponding PES does not show a gap with the correct quasi-hole counting (see Supplementary Note 4 and Supplementary Figure 5). This, however, should not be surprising: since the $\nu=2/5$ RR phase can be understood as the RR phase for holes at 3/5 hole-filling, adding holes to the system directly increases the ground-state energy as the 4-body interaction can no longer be minimized.
@@
Instead, the $\nu=2/5$ RR phase allows a certain number of quasi-particle excitations without increasing the ground-state energy. This is reflected in the hole entanglement spectrum (HES), where the many-body system is divided into $A$ and $B$ subsystems with $N_{A}$ and $N_{B}=N_{\text{s}}-N_{\text{e}}-N_{A}$ now denoting the number of holes. Importantly, the HES (Fig. 3(b)) displays an entanglement gap below which the number of states matches the quasi-particle counting of the 3/5 hole-filling (i.e. $\nu=2/5$) RR phase, indicating that the calculated states indeed correspond to the non-Abelian RR phase. From a topological perspective, this then implies that the nature of the RR states remains largely unaffected by the renormalization of the hole dispersion that is expected to arise from fluctuations in the quantum metric [72], even though these fluctuations may alter the ground state energies as suggested by the absence of a gap in Fig. 3(a).